# Stability of fully Discrete Local Discontinuous Galerkin method for the generalized Benjamin-Ono equation

### v5

@@ Abstract.
The main purpose of this paper is to design a fully discrete local discontinuous Galerkin (LDG) scheme for the generalized Benjamin-Ono equation. First, we analyze the stability for the semi-discrete LDG scheme and we prove that the scheme is $L^{2}$-stable for general nonlinear flux. We develop a fully discrete LDG scheme using the Crank-Nicolson (CN) method and fourth-order fourth-stage Runge-Kutta (RK) method in time. Adapting the methodology established for the semi-discrete scheme, we demonstrate the stability of the fully discrete CN-LDG scheme for general nonlinear flux. Additionally, we consider the fourth-order RK-LDG scheme for higher order convergence in time and prove that it is strongly stable under an appropriate time step constraint by establishing a three-step strong stability estimate for linear flux. Numerical examples are provided to validate the efficiency and optimal order of accuracy for both the methods.
@@
Local discontinuous Galerkin method, Benjamin-Ono equation, Hilbert transform, stability.
@@ 1. Introduction
The Benjamin-Ono equation (1.1) is a nonlinear, non-local partial differential equation that finds application in various physical phenomena [17]. In particular, the propagation of weakly nonlinear internal long waves in a fluid with a thin region of stratification can be represented by the Benjamin-Ono equation. Originating from the modeling of waves in shallow water, it offers insights into the behavior of these waves, including their propagation and interaction. Furthermore, we mention that it defines a Hamiltonian system, and with the help of inverse scattering method (see [1]), families of localized solitary wave solutions, called as soliton solutions [5], can be obtained under the appropriate assumptions on the initial data. Since the Benjamin-Ono equation is completely integrable, it admits infinitely many conserved quantities [5].
The investigation into the well-posedness of the Cauchy problem (1.1) associated to the Benjamin-Ono equation has been the subject of extensive research over the years. Pioneering work in the local well-posedness was conducted by Iório [16] for the initial data in $H^{s}(\mathbb{R})$, $s>3/2$ and making use of the conserved quantities, the global well-posedness for data in $H^{s}(\mathbb{R})$, $s\geq 2$ is demonstrated. Tao in [22] also obtained the global well-posedness in $H^{1}(\mathbb{R})$ by introducing the Gauge transformation. The idea of Gauge transformation given in [22] was further improved by Kenig et al. [15] to carry out the local well-posedness to $H^{s}(\mathbb{R})$ for $s\geq 0$. Molinet [19] has also obtained the global well-posedness for the periodic data in $L^{2}(\mathbb{R})$.
It is well-known that due to the effects of dispersion and nonlinear convection, finding a reliable method for the Benjamin-Ono equation is quite challenging task. Nevertheless, recent decades have seen the development of various numerical methods to solve the equation (1.1), with only the pertinent literature referenced here. An implicit finite difference method introduced by Thomée et al. [23] utilizes the continuous Hilbert transform. More recently, Dutta et al. [8] demonstrated the convergence of the fully discrete finite difference scheme, which includes the discrete Hilbert transform, and Galtung [12] devised a convergent Crank-Nicolson Galerkin scheme.
@@
The LDG method has been developed to deal with equations that have higher-order derivative terms. For instance, Yan and Shu [28] devised a LDG method for KdV type equations, which have third-order spatial derivatives. Furthermore, Xu and Shu [27] expanded the LDG method to handle equations with fourth and fifth order spatial derivatives.
In more recent times, the LDG method has become popular for dealing with partial differential equations that involve the non-local operator. Xu and Hesthaven [25] came up with an LDG method that breaks down the fractional Laplacian of order $\alpha$ $(1<\alpha<2)$ into second-order derivatives and fractional integrals of order $2-\alpha$. Similarly, Aboelenen [2] and Dwivedi et al. [11] developed a LDG method specifically designed for fractional Schrödinger-type equations and fractional Korteweg-de Vries equation respectively.
@@
In this paper, our approach to design the LDG scheme for the Benjamin-Ono equation (1.1) involves introduction of auxiliary variables to represent (1.1) into the system with lower order derivatives. A crucial part of this process involves constructing appropriate numerical fluxes at the interior interfaces, while boundary numerical fluxes are determined by the prescribed boundary conditions. Additionally, to develop a fully discrete LDG scheme, we discretize time using the Crank-Nicolson method. We demonstrate that the fully discrete scheme is stable for any general nonlinear flux. The main ingredients of the paper are enlisted below:
@@
    We design a LDG scheme for the Benjamin-Ono equation (1.1) and establish the stability of the devised semi-discrete scheme with a general nonlinear flux. The stability analysis has also been extended for fully discrete Crank-Nicolson scheme.
@@
The rest of the paper is organized as follows. We commence our investigation by introducing a few preliminary lemmas and semi-discrete LDG scheme and its stability analysis in Section 2. We present the stability analysis of the fully discrete Crank-Nicolson LDG scheme for general nonlinear flux and fully discrete fourth-oreder RK-LDG scheme for linear flux in Section 3. The efficiency of the scheme is validated through some numerical examples presented in Section 4. Concluding remarks and a few remarks about future work are given in Section 5.
@@ 2.1. Preliminary results
Hereby we describe a few relevant properties of the Hilbert transform through the following lemma. It is worthwhile to mention that these properties are instrumental for subsequent analysis. Let $\mathcal{S}(\mathbb{R})$ denotes the Schwartz space.
@@
(See [18, Chapter 15]) Let $\phi\in\mathcal{S}(\mathbb{R})$. Then the Hilbert transform $\mathcal{H}$ satisfies the following properties:
@@
Given that the original problem is defined over the entire real line due to the involvement of a non-local operator $\mathcal{H}$. However, for the numerical purpose, following the similar approach in [25, 23], we restrict it to a sufficiently large bounded domain $\Omega:=[a,b]$, where $a<b$, such that $U$ has a compact support within $\Omega$ for all time $0<t<T$. Hence it becomes imperative to impose the boundary conditions: $U(a,t)=0=U(b,t),\quad\text{for all }t<T$. Moreover, the properties of the Hilbert transform introduced in the Lemma 2.1 remain applicable for a bounded domain $\Omega$, provided $\phi$ has a compact support within $\Omega$.
@@
With all these preparation we introduce the auxiliary variables $P$ and $Q$ such that
@@
where $P^{k}(I_{i})$ is space of polynomials of degree up to order $k$ $(\geq 1)$ on $I_{i}$.
@@ 2.2. LDG scheme
Applying the DG approach in all the equations of the above system (2.2), we design the scheme as follows: we seek an approximations
@@
Hereby we analyze the stability the proposed semi-discrete LDG scheme (2.7) for (1.1).
@@ 2.3. Stability of semi-discrete scheme
($L^{2}$-stability) Let $u,p,q$ be obtained from the LDG scheme (2.7). Then the LDG scheme (2.7) for Benjamin-Ono equation (1.1) is $L^{2}$-stable. We have the following estimate
@@
Since we have the compact form (2.2) of the scheme (2.7), we choose the test functions $(v,w,z)=(u,-q,p)$ in (2.2), which yields
@@
for all $i=1,2,\cdots,N-1$. Dropping the positive term from left-hand side of equation (2.3) and using the boundary conditions, we end up with

### v6

@@ Abstract.
The main purpose of this paper is to design a fully discrete local discontinuous Galerkin (LDG) scheme for the generalized Benjamin-Ono equation. First, we proved the $L^{2}$-stability for the proposed semi-discrete LDG scheme and obtained a sub-optimal order of convergence for general nonlinear flux. We develop a fully discrete LDG scheme using the Crank-Nicolson (CN) method and fourth-order fourth-stage Runge-Kutta (RK) method in time. Adapting the methodology established for the semi-discrete scheme, we demonstrate the stability of the fully discrete CN-LDG scheme for general nonlinear flux. Additionally, we consider the fourth-order RK-LDG scheme for higher order convergence in time and prove that it is strongly stable under an appropriate time step constraint by establishing a three-step strong stability estimate for linear flux. Numerical examples associated with soliton solutions are provided to validate the efficiency and optimal order of accuracy for both methods.
@@
Local discontinuous Galerkin method, Benjamin-Ono equation, Hilbert transform, stability, error analysis.
@@ 1. Introduction
The Benjamin-Ono equation (1.1) is a nonlinear, non-local partial differential equation that finds application in various physical phenomena [18]. In particular, the propagation of weakly nonlinear internal long waves in a fluid with a thin region of stratification can be represented by the Benjamin-Ono equation. Originating from the modeling of waves in shallow water, it offers insights into the behavior of these waves, including their propagation and interaction. Furthermore, we mention that it defines a Hamiltonian system, and with the help of the inverse scattering method (see [1]), families of localized solitary wave solutions, called soliton solutions [5], can be obtained under the appropriate assumptions on the initial data. Since the Benjamin-Ono equation is completely integrable, it admits infinitely many conserved quantities [5].
The investigation into the well-posedness of the Cauchy problem (1.1) associated to the Benjamin-Ono equation has been the subject of extensive research over the years. Pioneering work in the local well-posedness was conducted by Iório [17] for the initial data in $H^{s}(\mathbb{R})$, $s>3/2$, and making use of the conserved quantities, the global well-posedness for data in $H^{s}(\mathbb{R})$, $s\geq 2$ is demonstrated. Tao in [24] also obtained the global well-posedness in $H^{1}(\mathbb{R})$ by introducing the gauge transformation. The idea of gauge transformation given in [24] was further improved by Kenig et al. [16] to carry out the local well-posedness to $H^{s}(\mathbb{R})$ for $s\geq 0$. Molinet [21] has also obtained the global well-posedness for the periodic data in $L^{2}(\mathbb{R})$.
It is well-known that due to the effects of dispersion and nonlinear convection, finding a reliable method for the Benjamin-Ono equation is quite a challenging task. Nevertheless, recent decades have seen the development of various numerical methods to solve the equation (1.1), with only the pertinent literature referenced here. An implicit finite difference method introduced by Thomée et al. [25] utilizes the continuous Hilbert transform. More recently, Dutta et al. [9] demonstrated the convergence of the fully discrete finite difference scheme, which includes the discrete Hilbert transform, and Galtung [13] devised a convergent Crank-Nicolson Galerkin scheme.
@@
The LDG method has been developed to deal with equations that have higher-order derivative terms. For instance, Yan and Shu [32] devised a LDG method for KdV type equations, which have third-order spatial derivatives. They obtained the error estimates with order of convergence $k+1/2$ in the linear case. Afterwards, Xu and Shu [29, 30] studied nonlinear convection-diffusion type equations and observed that the LDG method still provides similar levels of accuracy and order of convergence. Furthermore, Xu and Shu [31] expanded the LDG method to handle equations with fourth and fifth order spatial derivatives. Levy et al. [20] also worked on adapting the LDG method for equations with compactly supported traveling wave solutions appearing in nonlinear dispersive equations.
In more recent times, the LDG method has become popular for dealing with partial differential equations that involve the non-local operator. Xu and Hesthaven [27] came up with an LDG method that breaks down the fractional Laplacian of order $\alpha$ $(1<\alpha<2)$ into second-order derivatives and fractional integrals of order $2-\alpha$. This method turned out to be very effective, giving the optimal rates of $k+1$ in the linear case and $k+1/2$ in the nonlinear setup. Similarly, Aboelenen [2] and Dwivedi et al. [12] developed an LDG method specifically designed for fractional Schrödinger-type equations and fractional Korteweg-de Vries equations, respectively.
@@
In this paper, our approach to design the LDG scheme for the Benjamin-Ono equation (1.1) involves the introduction of auxiliary variables to represent (1.1) into the system with lower-order derivatives. A crucial part of this process involves constructing appropriate numerical fluxes at the interior interfaces, while boundary numerical fluxes are determined by the prescribed boundary conditions. Additionally, to develop a fully discrete LDG scheme, we discretize time using the Crank-Nicolson method. We demonstrate that the fully discrete scheme is stable for any general nonlinear flux. To the best of our knowledge, this is the first study to develop and analyze a DG method for the Benjamin-Ono equation. The proposed LDG scheme thus introduces a new framework for stable and efficient numerical approximations of this nonlocal dispersive equation. The main ingredients of the paper are enlisted below:
@@
    We design a LDG scheme for the Benjamin-Ono equation (1.1) and establish the stability and sub-optimal order of convergence of the devised semi-discrete scheme with a general nonlinear flux. The stability analysis has also been extended for fully discrete Crank-Nicolson scheme.
@@
The rest of the paper is organized as follows. We commence our investigation by introducing a few preliminary lemmas and a semi-discrete LDG scheme and its stability and error analysis in Section 2. We present the stability analysis of the fully discrete Crank-Nicolson LDG scheme for general nonlinear flux and the fully discrete fourth-order RK-LDG scheme for linear flux in Section 3. The efficiency of the scheme is validated through some numerical examples presented in Section 4. Concluding remarks and a few remarks about future work are given in Section 5.
@@ 2.1. Preliminary results
Hereby we describe a few relevant properties of the Hilbert transform through the following lemma. It is worthwhile to mention that these properties are instrumental for subsequent analysis. Let $\mathcal{S}(\mathbb{R})$ denote the Schwartz space.
@@
(See [19, Chapter 15]) Let $\phi\in\mathcal{S}(\mathbb{R})$. Then the Hilbert transform $\mathcal{H}$ satisfies the following properties:
@@
Given that the original problem is defined over the entire real line due to the involvement of a non-local operator $\mathcal{H}$. However, for the numerical purposes, following a similar approach in [27, 25], we restrict it to a sufficiently large bounded domain $\Omega:=[a,b]$, where $a<b$, such that $U$ has a compact support within $\Omega$ for all time $0<t<T$. Hence, it becomes imperative to impose the boundary conditions $U(a,t)=0=U(b,t),\text{ for all }t<T$. Moreover, the properties of the Hilbert transform introduced in the Lemma 2.1 remain applicable for a bounded domain $\Omega$, provided $\phi$ has a compact support within $\Omega$.
@@
With all this preparation we introduce the auxiliary variables $P$ and $Q$ such that
@@
where $P^{k}(I_{i})$ is the space of polynomials of degree up to order $k$ $(\geq 1)$ on $I_{i}$.
@@ 2.2. LDG scheme
Applying the DG approach in all the equations of the above system (2.2), we design the scheme as follows: we seek an approximation
@@
Hereby we analyze the stability of the proposed semi-discrete LDG scheme (2.7) for (1.1).
@@ 2.3. Stability of semi-discrete scheme
($L^{2}$-stability) Let $u,p,q$ be obtained from the LDG scheme (2.7). Then the LDG scheme (2.7) for the Benjamin-Ono equation (1.1) is $L^{2}$-stable. We have the following estimate
@@
Given the compact form (2.2) for the scheme (2.7), we choose the test functions $(v,w,z)=(u,-q,p)$ in (2.2), which yields
@@
for all $i=1,2,\cdots,N-1$. Dropping the positive term from the left-hand side of equation (2.3) and using the boundary conditions, we end up with
@@
2.4. Error analysis of the scheme
To advance with the error estimates, we define the special projection operators into the finite element space $V^{k}$ as follows. For any sufficiently smooth $g$, we define:
[…]
for all $i=1,2,\ldots,N$. Here $\mathcal{P}^{-}$ represents the projection defined above and $\mathcal{P}$ is the standard $L^{2}$ projection. Let $U$ be the exact solution to (1.1), and let $u$ be the approximate solution obtained through the LDG scheme (2.7). We introduce some compact notations for the terms involving the difference between projections and approximations by the following
$\mathcal{P}_{h}^{-}u=\mathcal{P}^{-}U-u,\quad\mathcal{P}_{h}q=\mathcal{P}Q-q,\quad\mathcal{P}_{h}p=\mathcal{P}P-p,$
and
$\mathcal{P}^{-}_{e}U=\mathcal{P}^{-}U-U,\quad\mathcal{P}_{e}Q=\mathcal{P}Q-Q,\quad\mathcal{P}_{e}P=\mathcal{P}P-P.$
Note that subscripts $h$ and $e$ in the above notations indicate the differences with approximations and exact solutions respectively.
In the process of establishing the error estimate for the equation (1.1), we introduce several lemmas concerning the relationship between the physical flux $f$ and the numerical flux $\hat{f}$.
Lemma 2.4 (See Lemma 3.1 in [33]).
Let $\xi\in L^{2}(\Omega)$ be any piecewise smooth function and $f$ be a nonzero $C^{1}$ flux. On each interface of elements and on the boundary points we define
$\beta(\hat{f};\xi)=\begin{cases}(f(\{\!\!\{\xi\}\!\!\})-\hat{f}(\xi))\llbracket\xi\rrbracket^{-1},\qquad&if~{}\llbracket\xi\rrbracket\neq 0,\\ \frac{1}{2}|f^{\prime}(\{\!\!\{\xi\}\!\!\})|,\qquad&if~{}\llbracket\xi\rrbracket=0,\end{cases}$
where $\hat{f}(\xi)$ is a consistent and monotone numerical flux. Then $\beta(\hat{f};\xi)$ is bounded and positive.
Borrowing the idea from [30], we look to estimate the nonlinear part $f(u)$ by defining
$\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,v)$ $=\sum\limits_{i=1}^{N}\int_{I_{i}}\big{(}f(U)-f(u)\big{)}v_{x}\,dx+\sum\limits_{i=1}^{N}\Big{(}\big{(}f(U)-f(\{\!\!\{u\}\!\!\})\big{)}\llbracket v\rrbracket\Big{)}_{i+\frac{1}{2}}$
$\quad+\sum\limits_{i=1}^{N}\Big{(}(f(\{\!\!\{u\}\!\!\})-\hat{f})\llbracket v\rrbracket\Big{)}_{i+\frac{1}{2}}.$ (2.22)
Lemma 2.5 (See Corollary 3.6 in [30]).
Let $f\in C^{3}(\Omega)$ and let the operator $\mathcal{G}_{i}$ be defined by (2.22). Then we have the following estimate:
$\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,v)$ $\leq-\frac{1}{4}\beta(\hat{f};u)\sum\limits_{i=1}^{N}\llbracket v\rrbracket^{2}_{i+\frac{1}{2}}+\big{(}C+C_{\ast}h^{-1}\left\|U-u\right\|_{L^{\infty}(\Omega)}^{2}\big{)}h^{2k+1}$
$\qquad+\big{(}C+C_{\ast}(\left\|v\right\|_{L^{\infty}(\Omega)}+h^{-1}\left\|U-u\right\|_{L^{\infty}(\Omega)}^{2})\big{)}\left\|v\right\|_{L^{2}(\Omega)}^{2}.$ (2.23)
We deal with the nonlinear flux $f(u)$ by making an $a~{}priori$ assumption [30]. Let $h$ be small enough and for $k\geq 1$, there holds
$\left\|U-u\right\|_{L^{2}(\Omega)}\leq h,$ (2.24)
where $u\in V^{k}$ is an approximation of $U$. The above assumption is unnecessary for linear flux $f(u)=u$. We define the bilinear operator $\mathcal{B}_{0}$ by the following
$\mathcal{B}_{0}(u,p,q;v,w,z)$ $=\sum\limits_{i=1}^{N}\Big{[}\left(u_{t},v\right)_{I_{i}}+\left(p,w\right)_{I_{i}}-\left(\mathcal{H}q,w\right)_{I_{i}}+\left(q,z\right)_{I_{i}}\Big{]}-\mathcal{D}^{+}(p,v)-\mathcal{D}^{-}(u,z)$
$=\sum\limits_{i=1}^{N}\Big{[}\left(u_{t},v\right)_{I_{i}}+\left(p,v_{x}\right)_{I_{i}}+\left(p,w\right)_{I_{i}}-\left(\mathcal{H}q,w\right)_{I_{i}}+\left(q,z\right)_{I_{i}}+\left(u,z_{x}\right)_{I_{i}}\Big{]}$
$\quad+\mathcal{IF}_{0}(u,p;v,z),$ (2.25)
where the term $\mathcal{IF}_{0}$ is given by
$\begin{split}\mathcal{IF}_{0}(u,p;v,z):=p^{+}_{\frac{1}{2}}v^{+}_{\frac{1}{2}}-p^{-}_{N+\frac{1}{2}}v^{-}_{N+\frac{1}{2}}+\sum\limits_{i=1}^{N-1}p^{+}_{i+\frac{1}{2}}\llbracket v\rrbracket_{i+\frac{1}{2}}+\sum\limits_{i=1}^{N-1}u^{-}_{i+\frac{1}{2}}\llbracket z\rrbracket_{i+\frac{1}{2}}.\end{split}$ (2.26)
Note that $\mathcal{B}_{0}=\mathcal{B}$ if we take $f=0$ in the definition of $\mathcal{B}$ in (2.2), that is, $\mathcal{B}_{0}$ is the linear part of $\mathcal{B}$. Incorporating (2.3) in (2.4), we have
$\begin{split}\mathcal{B}_{0}(u,p,q;u,-q,p)&=\left(u_{t},u\right)_{L^{2}(\Omega)}.\end{split}$ (2.27)
Theorem 2.6.
Let $U\in C^{k+1}(\Omega)$ be an exact solution of (1.1) and $u$ be an approximate solution obtained using the LDG scheme (2.7). Furthermore, assume that $f\in C^{3}(\Omega)$. Then, for sufficiently small $h$, the following error estimate holds
$\left\|U-u\right\|_{L^{2}(\Omega)}\leq Ch^{k+1/2},$ (2.28)
where $C$ is a constant depending on the fixed time $T>0$, $k\geq 1$, and the bounds on the derivatives $|f^{(m)}|$ for $m=1,2,3$.
Proof.
We begin by deriving an error equation. Since $U$ is an exact solution of (1.1), we define
$P=Q,\qquad Q=U_{x}.$
Then $U$, $P$, and $Q$ satisfy the equation (2.2). Hence for any $(v,w,z)\in V^{k}$, we have
$\mathcal{B}(U,P,Q;v,w,z)=\mathcal{B}(u,p,q;v,w,z)=0,$
where $(u,p,q)$ is an approximate solution obtained by the scheme (2.2). By incorporating the bilinear operator $\mathcal{B}_{0}$, we get
00 0 $=\mathcal{B}(U,P,Q;v,w,z)-\mathcal{B}(u,p,q;v,w,z)$
$=\mathcal{B}_{0}(U,P,Q;v,w,z)-\mathcal{B}_{0}(u,p,q;v,w,z)-\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,v)$
$=\mathcal{B}_{0}(U-u,P-p,Q-q;v,w,z)-\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,v).$
Taking into account the projection operators $\mathcal{P}^{-}$ and $\mathcal{P}$ defined in (2.21), we choose $(v,w,z)=(\mathcal{P}_{h}^{-}u,-\mathcal{P}_{h}q,\mathcal{P}_{h}p)$. Since
$U-u=\mathcal{P}^{-}_{h}u-\mathcal{P}^{-}_{e}U,\quad P-p=\mathcal{P}_{h}p-\mathcal{P}_{e}P\quad\text{ and }\quad Q-q=\mathcal{P}_{h}q-\mathcal{P}_{e}Q,$
we have
$\begin{split}\mathcal{B}_{0}(\mathcal{P}_{h}^{-}u,\mathcal{P}_{h}p,\mathcal{P}_{h}q;\mathcal{P}_{h}^{-}u,-\mathcal{P}_{h}q,\mathcal{P}_{h}p)&=\mathcal{B}_{0}(\mathcal{P}^{-}_{e}U,\mathcal{P}_{e}P,\mathcal{P}_{e}Q;\mathcal{P}_{h}^{-}u,-\mathcal{P}_{h}q,\mathcal{P}_{h}p)\\ &\quad+\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,\mathcal{P}_{h}^{-}u).\end{split}$ (2.29)
We estimate the first term on the right-hand side of (2.29). From equation (2.4), we have
$\begin{split}\mathcal{B}_{0}(\mathcal{P}^{-}_{e}U,\mathcal{P}_{e}P,\mathcal{P}_{e}Q;\mathcal{P}_{h}^{-}u,-\mathcal{P}_{h}q,\mathcal{P}_{h}p)&=\sum\limits_{i=1}^{N}\Big{[}\left((\mathcal{P}^{-}_{e}U)_{t},\mathcal{P}_{h}^{-}u\right)_{I_{i}}+\left(\mathcal{P}_{e}P,(\mathcal{P}_{h}^{-}u)_{x}\right)_{I_{i}}\\ &\quad-\left(\mathcal{P}_{e}P,\mathcal{P}_{h}q\right)_{I_{i}}+\left(\mathcal{H}\mathcal{P}_{e}Q,\mathcal{P}_{h}q\right)_{I_{i}}\\ &\quad+\left(\mathcal{P}_{e}Q,\mathcal{P}_{h}p\right)_{I_{i}}+\left(\mathcal{P}^{-}_{e}U,(\mathcal{P}_{h}p)_{x}\right)_{I_{i}}\Big{]}\\ &\quad+\mathcal{IF}_{0}(\mathcal{P}^{-}_{e}U,\mathcal{P}_{e}P;\mathcal{P}_{h}^{-}u,\mathcal{P}_{h}p).\end{split}$ (2.30)
Since we have
$(\mathcal{P}_{h}p)_{x}\in P^{k-1}(I_{i}),~{}(\mathcal{P}_{h}^{-}u)_{x}\in P^{k-1}(I_{i}),\quad\text{and}\quad\mathcal{P}_{h}p,~{}\mathcal{P}_{h}q\in P^{k}(I_{i}),$
and consequently, from the projection properties defined in (2.21) implies
$\left(\mathcal{P}_{e}P,(\mathcal{P}_{h}^{-}u)_{x}\right)_{I_{i}}=0,\quad\left(\mathcal{P}_{e}Q,\mathcal{P}_{h}p\right)_{I_{i}}=0,\quad\left(\mathcal{P}^{-}_{e}U,(\mathcal{P}_{h}p)_{x}\right)_{I_{i}}=0,$
for all $i=1,2,\cdots,N$, and $(\mathcal{P}^{-}_{e}U)^{-}_{i+\frac{1}{2}}=0$ for all $i=1,2,\cdots,N-1$. Furthermore, we observe that
$\sum\limits_{i=1}^{N}\left(\mathcal{P}_{e}P-\mathcal{H}\mathcal{P}_{e}Q,\mathcal{P}_{h}q\right)_{I_{i}}\leq\|(\mathcal{P}P-P)-\mathcal{H}(\mathcal{P}Q-Q)\|_{L^{2}(\Omega)}\|\mathcal{P}_{h}q\|_{L^{2}(\Omega)}\leq C(\varepsilon)h^{2k+2}+\varepsilon\|\mathcal{P}_{h}q\|_{L^{2}(\Omega)}^{2},$ (2.31)
where we have used the isometry of the Hilbert transform from Lemma 2.1 and the projection property, and the Young’s inequality. Moreover, from the approximation theory on the point values associated with the projection operators [6, Section 3.2], we have
$\left(\mathcal{P}_{e}P\right)^{+}_{i+\frac{1}{2}}\leq Ch^{k+1},\quad(\mathcal{P}^{-}_{e}U)^{-}_{i-\frac{1}{2}}\leq Ch^{k+1},$
for all $i=1,2,\cdots,N-1$, and
$(\mathcal{P}_{h}^{-}u)^{-}_{N+\frac{1}{2}}=0,\quad(\mathcal{P}_{h}^{-}u)^{+}_{\frac{1}{2}}=0,$
by using boundary conditions. Combining these error bounds and using the Young’s inequality, we have
$\mathcal{IF}_{0}(\mathcal{P}^{-}_{e}U,\mathcal{P}_{e}P;\mathcal{P}_{h}^{-}u,\mathcal{P}_{h}p)=$ $(\mathcal{P}_{e}P)^{+}_{\frac{1}{2}}(\mathcal{P}_{h}^{-}u)^{+}_{\frac{1}{2}}-(\mathcal{P}_{e}P)^{-}_{N+\frac{1}{2}}(\mathcal{P}_{h}^{-}u)^{-}_{N+\frac{1}{2}}$
$+\sum\limits_{i=1}^{N-1}(\mathcal{P}_{e}P)^{+}_{i+\frac{1}{2}}\llbracket\mathcal{P}_{h}^{-}u\rrbracket_{i+\frac{1}{2}}+\sum\limits_{i=1}^{N-1}(\mathcal{P}^{-}_{e}U)^{-}_{i+\frac{1}{2}}\llbracket\mathcal{P}_{h}p\rrbracket_{i+\frac{1}{2}}$
$\leq$ $\sum\limits_{i=1}^{N-1}\Big{(}C(\varepsilon)\big{(}(\mathcal{P}_{e}P)^{+}_{i+\frac{1}{2}}\big{)}^{2}+\varepsilon\llbracket\mathcal{P}_{h}^{-}u\rrbracket_{i+\frac{1}{2}}^{2}\Big{)}$
$\leq$ $C(\Omega)h^{2k+1}+\varepsilon\sum\limits_{i=1}^{N-1}\llbracket\mathcal{P}_{h}^{-}u\rrbracket_{i+\frac{1}{2}}^{2},$
where $\varepsilon>0$. Using the above estimates in (2.30), we end up with
$\begin{split}\mathcal{B}_{0}(\mathcal{P}^{-}_{e}U,\mathcal{P}_{e}P,\mathcal{P}_{e}Q;\mathcal{P}_{h}^{-}u,-\mathcal{P}_{h}q,\mathcal{P}_{h}p)&\leq\left((\mathcal{P}^{-}_{e}U)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}+C(\Omega)h^{2k+1}+\varepsilon\sum\limits_{i=1}^{N-1}\llbracket\mathcal{P}_{h}^{-}u\rrbracket_{i+\frac{1}{2}}^{2}.\end{split}$ (2.32)
Incorporating estimates from (2.27) and (2.32) in (2.29), we have
$\left((\mathcal{P}_{h}^{-}u)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}\leq$ $\left((\mathcal{P}^{-}_{e}U)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}+C(\Omega)h^{2k+1}+\varepsilon\sum\limits_{i=1}^{N-1}\llbracket\mathcal{P}_{h}^{-}u\rrbracket_{i+\frac{1}{2}}^{2}$
$\quad+\varepsilon\|\mathcal{P}_{h}q\|_{L^{2}(\Omega)}^{2}+\sum\limits_{i=1}^{N}\mathcal{G}_{i}(f;U,u,\mathcal{P}_{h}^{-}u).$ (2.33)
With the help of estimate (2.5) and as $\varepsilon$ is arbitrarily small, the estimate (2.4) reduces to
$\begin{split}&\left((\mathcal{P}_{h}^{-}u)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}+\frac{1}{4}\beta(\hat{f};\mathcal{P}_{h}^{-}u)\sum\limits_{i=1}^{N}\llbracket\mathcal{P}_{h}^{-}u\rrbracket^{2}_{i+\frac{1}{2}}\\ &\qquad\leq\left((\mathcal{P}^{-}_{e}U)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}+C(\Omega)h^{2k+1}+\Big{(}C+C_{\ast}h^{-1}\left\|U-u\right\|_{L^{\infty}(\Omega)}^{2}\Big{)}h^{2k+1}\\ &\qquad\qquad+\Big{(}C+C_{\ast}\big{(}\left\|\mathcal{P}_{h}^{-}u\right\|_{L^{\infty}(\Omega)}+h^{-1}\left\|U-u\right\|_{L^{\infty}(\Omega)}^{2}\big{)}\Big{)}\left\|\mathcal{P}_{h}^{-}u\right\|_{L^{2}(\Omega)}^{2}.\end{split}$ (2.34)
Utilizing the inverse inequality $\left\|u\right\|_{L^{\infty}(\Omega)}\leq h^{-1/2}\left\|u\right\|_{L^{2}(\Omega)}$ and a priori assumption (2.24), we obtain the estimate
$h^{-1}\left\|U-u\right\|_{L^{\infty}(\Omega)}^{2}h^{2k+1}\leq h^{-2}\left\|U-u\right\|^{2}_{L^{2}(\Omega)}h^{2k+1}\leq h^{2k+1}.$ (2.35)
Using the above estimate and the positivity of $\beta$ from Lemma 2.4, equation (2.34) implies
$\frac{1}{2}\frac{d}{dt}\left\|\mathcal{P}_{h}^{-}u\right\|^{2}_{L^{2}(\Omega)}\leq\left((\mathcal{P}^{-}_{e}U)_{t},\mathcal{P}_{h}^{-}u\right)_{L^{2}(\Omega)}+C({\Omega})h^{2k+1}+C\left\|\mathcal{P}_{h}^{-}u\right\|^{2}_{L^{2}(\Omega)}.$
Using the standard approximation theory associated to the projection [6, Section 3.2], we have $\left\|\mathcal{P}_{h}^{-}u(\cdot,0)\right\|^{2}_{L^{2}(\Omega)}=0$. Finally, with the help of Gronwall’s inequality, we obtain the estimate
$\left\|U-u\right\|_{L^{2}(\Omega)}\leq Ch^{k+1/2}.$
This completes the proof. ∎